# A dynamic view of some anomalous phenomena in SGD

### v2

@@ 1 Introduction
    For a deterministic or random sequence $\{w_{n}\}$, $w_{n}=O(n)$ will stand for $\limsup_{n\to\infty}\frac{|w_{n}|}{n}<\infty$ (‘a.s.’ in the random case). If $w_{n}\neq 0$, we denote by $w_{n}=\Theta(n)$ the condition $w_{n}=O(n)$ and $w^{-1}_{n}=O(n)$ simultaneously (‘a.s.’ in the random case). For vector-valued $\{w_{n}\}$, $w_{n}=O(n)$ or $\Theta(n)$ if it is so componentwise.
@@ 2 A theorem of Tim Austin
Corollary 2.2.
Let $D:=$ a compact convex set in $\mathbb{R}^{d},d\geq 1,$ with nonempty interior and smooth boundary, and let $f$ above be continuously differentiable. Then we can replace the $L_{1}$-norm in (2.1) by the sup norm.
Proof.
This is an immediate consequence of the Sobolev embedding theorem [3]. Specifically, let $W^{1,p}(D):=$ the space of functions differentiable in the sense of distributions for which the $p$-th moment, $p\geq 1$, of the function and its distributional partial derivatives is bounded. It is endowed with the norm
$\|f\|_{1,p}:=\left(\int|f|^{p}d\mu\right)^{\frac{1}{p}}+\sum_{i}\left(\int\left|\frac{\partial f}{\partial x_{i}}\right|^{p}d\mu\right)^{\frac{1}{p}},\ f\in W^{1,p}(D).$
When $p>N$, this space is continuously embedded in $C(D)$ equipped with the max norm $\|\ \cdot\ \|$, whence
$\|f\|_{1,p}\leq c\|f\|$
for some constant $c>0$. Specializing these observations to continuously differentiable functions, which are contained in $W^{1,p}(D)$ for all $p\geq 1$, the claim follows for this class of functions. ∎
This encapsulates the well known observation that a function of a very large number of variables will typically depends predominantly on a significantly smaller subset thereof. In view of these observations, our starting point will be a stylized model for stochastic gradient descent (SGD) in high dimensions that one encounters while training large over-parametrized neural networks. Consider a function $F:\mathbb{R}^{N}\to\mathbb{R}$ for some $N\gg 1$, which will serve as our surrogate for the map that maps the parameters to the expected loss for an over-parametrized neural network. Since our interest is in analyzing SGD for this function, we shall assume that it is continuously differentiable.
@@ 4 Two time scale iterations
The above observation facilitates analysis of the SGD in (3.3)-(3.4) as a two time scale stochastic approximation, with the difference (as opposed to [12]) being that we have constant step sizes $a\gg b>0$. In the classical theory of [12], $a,b$ are replaced by decreasing step sizes $a(n),b(n),n\geq 0,$ satisfying the Robbins-Monro conditions $\sum_{n}a(n)=\sum_{n}b(n)=\infty$ and $\sum a(n)^{2},\sum_{n}b(n)^{2}<\infty$, along with the additional condition $b(n)=o(a(n))$ as $n\to\infty$. Treating $a(n)$ as a discrete time step, define continuous piecewise linear interpolations of $\{x(n)\},\{y(n)\}$ as follows : Let $t(0)=0$, $t(n+1)=t_{n}+a(n)$ for $n\geq 0$. Set $\bar{x}(t(n))=x(n),\bar{y}(t(n))=y(n)$ with linear interpolation on $t(n),t(n+1)],n\geq 0$. As argued in [12] (also [14], Section 8.1), one can show that
@@ 5 The anomalous behaviour
    The the above statements amount to the claims from [12], [14] specialized to SGD.
@@ 7 Conclusions
    The small noise limit argument in the last section is handicapped by the fact that very little literature is available regarding this, unlike the single time scale case. There is ample opportunity to develop the basic theory itself much further and then apply it to algorithmic issues such as those considered here.

### v3

@@ 1 Introduction
    For a deterministic or random sequence $\{w_{n}\}$, $w_{n}=O(n)$ will stand for $\limsup_{n\to\infty}\frac{|w_{n}|}{n}<\infty$ (‘a.s.’ in the random case). If $w_{n}\neq 0$, we denote by $w_{n}=\Theta(n)$ the conditions $w_{n}=O(n)$ and $w^{-1}_{n}=O(n)$ simultaneously (‘a.s.’ in the random case). For vector-valued $\{w_{n}\}$, $w_{n}=O(n)$ or $\Theta(n)$ if it is so componentwise.
@@ 2 A theorem of Tim Austin
This encapsulates the well known observation that a function of a very large number of variables will typically depend predominantly on a significantly smaller subset thereof. In view of these observations, our starting point will be a stylized model for stochastic gradient descent (SGD) in high dimensions that one encounters while training large over-parametrized neural networks. Consider a function $F:\mathbb{R}^{N}\to\mathbb{R}$ for some $N\gg 1$, which will serve as our surrogate for the map that maps the parameters to the expected loss for an over-parametrized neural network. Since our interest is in analyzing SGD for this function, we shall assume that it is continuously differentiable.
@@ 4 Two time scale iterations
The above observation facilitates analysis of the SGD in (3.3)-(3.4) as a two time scale stochastic approximation, with the difference (as opposed to [12]) being that we have constant step sizes $a\gg b>0$. In the classical theory of [12], $a,b$ are replaced by decreasing step sizes $a(n),b(n),n\geq 0,$ satisfying the Robbins-Monro conditions $\sum_{n}a(n)=\sum_{n}b(n)=\infty$ and $\sum a(n)^{2},\sum_{n}b(n)^{2}<\infty$, along with the additional condition $b(n)=o(a(n))$ as $n\to\infty$. Treating $a(n)$ as a discrete time step, define continuous piecewise linear interpolations of $\{x(n)\},\{y(n)\}$ as follows : Let $t(0)=0$, $t(n+1)=t(n)+a(n)$ for $n\geq 0$. Set $\bar{x}(t(n))=x(n),\bar{y}(t(n))=y(n)$ with linear interpolation on $t(n),t(n+1)],n\geq 0$. As argued in [12] (also [14], Section 8.1), one can show that
@@ 5 The anomalous behaviour
    The above statements amount to the claims from [12], [14] specialized to SGD.
@@ 7 Conclusions
    The small noise limit argument in the last section is handicapped by the fact that very little literature is available regarding this for two time scale iterations, unlike the single time scale case. There is ample opportunity to develop the basic theory itself much further and then apply it to algorithmic issues such as those considered here.